Formulate the situation as a system of two linear equations in two variables. Be sure to state clearly the meaning of your x- and y-variables. Solve the system by the elimination method. Be sure to state your final answer in terms of the original question.
A lawyer has found 60 investors for a limited partnership to purchase an inner-city apartment building, with each contributing either $3,000 or $6,000. If the partnership raised $246,000, then how many investors contributed $3,000 and how many contributed $6,000?
x = $3,000 investors
y = $6,000 investors

Answer by mananth(16946)


$6,000 x Investors 60
$3,000 y Investors $246,000
Equation for total investors

x + y = 60 .............1

Equation for amount raised
$6,000 x + 3000.00 y = $246,000 .............2
Eliminate y
multiply (1)by -3000
Multiply (2) by 1
-3000 x -3000 y = -180000
6000 3000 y = 246000
Add the two equations
3000 x = 66000
/ 3000.00
x = 22.00
plug value of x in (1)
1.00 x + 1.00 y = 60
22.00 + 1.00 y = 60
1.00 y = 38
y= 38

$6000.00 22.00 Investors
$3000.00 38.00 Investors
m.ananth@hotmail.ca